OpenStudy (anonymous):

An experiment consist of rolling 3 fairl dice--1 red die, a blue die, and a white die---and recording the number rolled on each die. Assume that the dice are fair, so that all outcomes are equally likely. 1. What probability should be assigned to each outcome?____ 2. What is the probability that the sum of the numbers rolled is 6? 3. What is the probability that the sum of the numbers rolled is at most 5?

OpenStudy (anonymous):

7/216 is the probability for rolling a 6. 10/216 or 5/108 is the probability of rolling 5 or less.

Directrix (directrix):

The possible outcomes for rolling 3-dice – organized according to the sum of the 3 dice. Click the link. http://homepage.smc.edu/mcgraw_colleen/math_52/dice%20roulette.pdf

Directrix (directrix):

2. What is the probability that the sum of the numbers rolled is 6? 10 ways to get a "six" sum out of 216 sums. P(6) = 10/ 216 = 5/108

Directrix (directrix):

What is the probability that the sum of the numbers rolled is at most 5? 10/216 or 5 / 108

Directrix (directrix):

1. What probability should be assigned to each outcome?____ In this instance, I don't know what is defined to be an outcome. The results of tossing all 3 die or the results of 1 die. All 216 outcomes are listed on the http://homepage.smc.edu/mcgraw_colleen/math_52/dice%20roulette.pdf

Directrix (directrix):

On #1 the probability of an outcome on tossing all 3 die would be 1/6 times 1/6 times 1/6 or 1/216.

Directrix (directrix):

CW, did you look at these?

Directrix (directrix):

7/216 is the probability for rolling a 6. NOT Corrrect

Directrix (directrix):

114, 141, 411, 123, 132, 213, 231, 321, 312, 222 for a total of ten that sum to 6.

OpenStudy (anonymous):

sorry disconnected and i signed off It was correct....thank you so much

Directrix (directrix):

yw